# Physics-Informed Detection of Friction Anomalies in Satellite Reaction Wheels

### v1

@@ V Results
The algorithm’s structure, composed of four sequential stages, is beneficial for analysis, as each part of the algorithm can be analyzed independently to ensure it works as expected. This way, our analysis is not limited to the final accuracy obtained by the last stage. Unfortunately, while we have a ground truth in the classification because the dataset is labeled, we do not have such a reference for the output of the different stages of the algorithm. Therefore, we will use various methods to assess other aspects of the work.
@@
Rather than organizing this section by objectives, we organize it by stages of the algorithm. We start by analyzing the changepoint detection stage using synthetic data. For the friction estimation and changepoint assignment, we use the TAS dataset to analyze some performance metrics so, even without a ground truth to compare to, we can study the behavior of these stages. Finally, we will use the labels provided in the TAS dataset to assess the accuracy of the complete algorithm at determining the anomaly present in the RWA. Additionally, we investigate the computational cost of the whole process to establish its suitability for running in a low-power device located in the satellite.
V-A Changepoint detection
@@
In the analysis of a changepoint detection algorithm, there are two main metrics of performance: the Missed Detection Rate (MDR), defined as the probability of not detecting a changepoint, and the Average Run Length (ARL), defined as the average number of points the algorithm goes through before detecting a false changepoint in the data [44]. Since it is always possible to improve one at the expense of the other (by decreasing or increasing the threshold $\text{GLR}_{\text{thr}}$), the challenge lies in achieving a high ARL with a low MDL.
To assess the performance of the changepoint detection, it is important to remember how the subsequent stages of the algorithm process changepoints. As mentioned in Section IV-C2, an undetected changepoint can invalidate the algorithm’s output while a false positive can be rejected during the assignment stage. This implies that while high ARL is beneficial, as it will lead to fewer false positives, a low MDR is fundamental for the algorithm’s success.
For our analysis, we will fix the MDR to $0.1\%$ and determine the threshold value at which such sensitivity is achieved using the $\chi^{2}$ distribution. Then, we obtain the average run length with this threshold, which we call the $\text{ARL}^{0.1\%}$. A satisfactory value of ARL depends on the tolerance of the algorithm to false positives, and the number of points we process in a single window of measurements. In the TAS case, we analyze windows of around $80000$ measurements, so an ARL of around $10^{6}$ is a good target if we do not want to get more than $1$ false changepoint per window. Since the TAS dataset has no annotated changepoint locations, these results are obtained using our own simulations, in which we can place changepoints in a controlled way.
V-A1 Bias selection
First, we analyze the effect of bias on the algorithm’s performance. Recalling Subsection IV-A, $W_{\mathrm{b}}$ is a tunable parameter, representing the weight of the prior on the viscous friction coefficient in (10). Since there is no clear way of choosing this value to obtain better results, we run a parametric study to observe how it affects the $\text{ARL}^{0.1\%}$ for different window sizes, assuming that the changes of dry friction we are trying to detect have a magnitude of $3$ times the noise standard deviation. The results are shown in Figure 7, left plot. A higher window size $w$ leads to better algorithm results, as the effect of the noise is reduced when we use more points the w-GLR. However, increasing $w$ is not always feasible, because it is limited by the minimum distance we expect to see between changepoints. We can observe how the $W_{\mathrm{b}}$ improves the algorithm’s performance for different window sizes. As an example, using a window size of $20$, we can improve the $ARL^{0.01\%}$ from $10^{4}$ to $10^{10}$ by increasing $W_{\mathrm{b}}$ from $10^{-8}$ to $10^{-4}$. The bias term is, then, a good way of improving the performance of the changepoint detection when the window size is constrained, and it seems from this first analysis that $W_{\mathrm{b}}$ should be set to be arbitrarily large.
However, in the previous plot, we assumed that our estimated friction $\tilde{f}^{\mathrm{v}}$ is equal to the actual viscous friction $f^{\mathrm{v}}$. Unfortunately, this may not be the case, since the latter is not known in advance. Considering the effect of the bias $\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}}$ is more complex: when the error is not zero, the w-GLR no longer follows a $\chi^{2}$ distribution. We can, however, estimate the effect of this error to be in the order of $W_{\mathrm{b}}(\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}})^{2}$. As the original distribution of the w-GLR is of order unit, we should ensure that $W_{\mathrm{b}}(\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}})^{2}\lll 1$ so the error does not have a noticeable effect. Using $W_{\mathrm{b}}=10^{-4}$, this implies that we need $|\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}}|\lll 10^{2}$, an easily achievable condition given that the viscous friction coefficient is of order $1$.
This gives a simple rule for tuning $W_{b}$: if the window size is large enough (more than 50 points), it can be set to 0. If not, a high value can be set to improve the results, but keeping in mind that if the prior of the viscous friction is not reliable, the performance may degrade and produce too many false positives.
V-A2 Changepoint properties
@@
We now analyze the effect of the window size $w$ and the size of the changepoints $\Delta f_{\mathrm{cp}}$ on the $\text{ARL}^{0.1\%}$, while keeping $W_{b}=10^{-4}$. The results are shown in Figure 7, right plot. We can see that better results are achieved with either high window sizes or changepoint strengths. This shows a basic limitation of the algorithm: if we want to detect small friction changes, we need to use large window sizes. However, this might not be possible if the changepoints that we want to detect may be too close together, in which case we may need to increase the value of $W_{\mathrm{b}}$.
@@
As a last note, the average run lengths obtained in these analyses assume that the data contains no abnormalities, that is, the friction is just a combination of dry and viscous friction and Gaussian noise as given by (4a). Real data may contain imperfections that trigger additional, unavoidable false changepoints. Therefore, a high ARL is necessary for reducing the number of false changepoints, but it does not guarantee it.
V-B Friction Estimation
@@
We now analyze the output of the second step of the algorithm: the friction estimation.
This algorithm step is of no significant interest since the Least-Squares algorithm has well-known properties, and the only parameter to adjust is $\Delta k$, which does not have a large impact on the output. We can, on the other hand, analyze how well the friction coefficients $\hat{f}^{\mathrm{d}}_{i}$ and $\hat{f}^{v}$ obtained in this stage approximate the friction $\hat{f}_{k}$ in the data.
@@
A small error is an indicator of a good fit of the model to the data (at least when it comes to (4a), since we are not considering the different FSM yet) and a correct estimation of the changepoint locations. This is because, if changes in the friction were not detected, the algorithm would try to fit the same coefficient to intervals of measurements with different dry friction coefficients, which would produce a high error.
@@
As in the previous subsection, there is a word of caution. Although a low RMSE is a good indicator of the correct fit of the coefficients, it is not a flawless metric. An algorithm that detects more changepoints than necessary will produce a reduced RMSE. Although the previous subsection showed that the false positive detection was not a cause of concern, the following subsection will explore the issue of false positives in more detail.
V-C Assignment
@@
N rejections $\mathbf{1}$ $\mathbf{2}$ $\mathbf{3}$ $\mathbf{4}$ $\mathbf{5}$ $\mathbf{6}$ $\boldsymbol{\geq}\mathbf{7}$
@@
Figure 10: Relation between the number of iterations run by the assignment algorithm before obtaining the maximum likelihood assignment, and the number of changepoints to assign. The mean of $116.6$ iterations was obtained across all datapoints in the TAS dataset.
@@
As a final note, we must remember that, on average, the number of iterations is relatively small, around 115 iterations. This is because, although this number can reach high values when the number of changepoints and/or rejections is high, both numbers are relatively small for most runs.
V-D Classification
@@
V-D1 Anomaly classification
To determine how accurate the algorithm is at determining the anomaly present in an RWA, we compare its output to the labels in the TAS dataset. Although the algorithm can determine several anomalies at once, assigning a binary label to each one, time-series in the dataset can only have one anomaly at most. We consider the five possible actual statuses (1 nominal + 4 anomalous), according to the label, and determine in what percentage of the data points with that status the algorithm detected each anomaly.
@@ V-D1 Anomaly classification
For the four anomalies considered, the probability of them being correctly detected is around $95\%$. The only exception is the FSS 2, which has an accuracy of $100\%$. Perfect accuracy in the classification implies that a separating plane exists between the histograms of the friction values of the FSS 2 with and without the anomaly.
@@
Finally, we observe that the differences between the lowest and highest values obtained in different trainings are very small, with the most significant variation around $5\%$. This can be attributed to the simplicity of the SVM, since a small number of tunable hyperparameters reduces the risk of overfitting.
V-D2 Number of bins
For the TAS dataset, we used 40 bins to generate the histograms used to classify anomalies of both switching systems. To support this decision, we tested different bin numbers ($n_{\mathrm{bins}}$) and computed the resulting train and validation loss in Figure 12.
The plot shows an expected decrease in the train loss as the number of bins increases, since the dimension of the classification increases with it. However, many bins lead to overfitting, as training captures only the specific shape of the datapoints used in training, and the validation loss increases. This shows a good level of robustness of the algorithm to the number of bins, which can again be attributed to the use of SVMs for the classification.
V-E Computation Time
We now evaluate the times it takes the algorithm to run on the datapoints of the dataset provided by TAS, each of them containing around 80000 measurements of spin rate and friction. The tests were run on a laptop with an Intel® Core™ Ultra 7 165U.

### v2

@@ VI Experimental Results
We now proceed to analyze the performance of the anomaly detection algorithm developed in this paper. The algorithm’s structure, composed of four sequential stages, is beneficial for analysis, as each part of the algorithm can be analyzed independently to ensure it works as expected. This way, our analysis is not limited to the final accuracy obtained by the last stage.
@@
This section is organized by stages of the algorithm. We start by analyzing the changepoint detection stage using synthetic data. For the friction estimation and changepoint assignment, we use the TAS dataset to analyze some performance metrics so that, even without a ground truth to compare to, we can study the behavior of these stages. Finally, we will use the labels provided in the TAS dataset to assess the accuracy of the complete algorithm at determining the anomaly present in the RWA. Additionally, we investigate the computational cost of the whole process to establish its suitability for running in a low-power device located in the satellite.
VI-A Changepoint detection
@@
The capability of the changepoint detector to locate actual changepoints without being triggered with false positives depends, given a fixed $W_{b}$, on elements: the window size $w$ and the change of friction $\Delta f_{\mathrm{cp}}$ at the changepoints. We use the same metric defined in Section V-1 to perform this analysis, studying the ARL obtained when the threshold of the detector is set to achieve a MDR of $0.1\%$, which we call the $\text{ARL}^{0.1\%}$. As with the tuning, we generate data artificially so we have control over where the actual changepoints are located.
@@
The results are shown in Figure 7, right plot. We can see that better results are achieved with either high window sizes or changepoint strengths. This shows a basic limitation of the algorithm: if we want to detect small friction changes, we need to use large window sizes. However, this might not be possible if the changepoints that we want to detect may be too close together, in which case we need to increase the value of $W_{\mathrm{b}}$.
@@
As a last note, the average run lengths obtained in this analysis assume that the data contains no abnormalities, that is, the friction is just a combination of dry and viscous friction and Gaussian noise as given by (4). Real data may contain imperfections that trigger additional, unavoidable false changepoints. Therefore, a high ARL is necessary for reducing the number of false changepoints, but it does not guarantee it.
VI-B Friction Estimation
@@
We now analyze the output of the second step of the algorithm: the friction estimation. In particular, we will analyze how well the friction coefficients $\hat{f}^{\mathrm{d}}_{i}$ and $\hat{f}^{\mathrm{v}}$ obtained in this stage approximate the friction $\hat{f}_{k}$ in the data.
@@
A small error is an indicator of a good fit of the model to the data (at least when it comes to (4), since we are not considering the different FSS yet) and a correct estimation of the changepoint locations. This is because, if changes in the friction were not detected, the algorithm would try to fit the same coefficient to intervals of measurements with different dry friction coefficients, which would produce a high error.
@@
As in the previous subsection, there is a word of caution. Although a low RMSE is a good indicator of the correct fit of the coefficients, it is not a flawless metric. An algorithm that detects more changepoints than necessary will produce a reduced RMSE. In the following subsection we will explore the issue of false positives in more detail.
VI-C Maximum Likelihood Assignment
@@
#\# rejected $\mathbf{1}$ $\mathbf{2}$ $\mathbf{3}$ $\mathbf{4}$ $\mathbf{5}$ $\mathbf{6}$ $\boldsymbol{\geq}\mathbf{7}$
@@
Figure 11: Relation between the number of iterations run by the assignment algorithm before obtaining the maximum likelihood assignment, and the number of changepoints to assign. A mean of $116.6$ iterations was obtained across all datapoints in the TAS dataset.
@@
As a final note, we must keep in mind that, on average, the number of iterations is relatively small, around 115 iterations. This is because, although this number can reach high values when the number of changepoints and/or rejections is high, both numbers are relatively small for most runs.
VI-D Anomaly Classification
@@
To determine the algorithm’s accuracy at determining the anomaly present in an RWA, we compare its output to the labels in the TAS dataset. Although the algorithm can determine several anomalies at once, time-series in the dataset only have one anomaly at most. We consider the five possible labeled statuses (nominal + 4 anomalies) and determine, for each, how likely each anomaly is to be detected by the algorithm.
@@
For the four anomalies considered, the probability of them being correctly detected is around $95\%$. The only exception is the FSS 2, which has an accuracy of $100\%$, which implies that a separating plane exists between the histograms of the friction values of the FSS 2 with and without the anomaly.
@@
Finally, we observe that the differences between the lowest and highest values obtained in different trainings are very small, with the most significant variation around $5\%$. This can be attributed to the simplicity of the SVM, as a small number of tunable hyperparameters reduces the risk of overfitting.
VI-E Computation Time
We now evaluate the times it takes the algorithm to run on the datapoints of the dataset provided by TAS, each containing around 80000 measurements of spin rate and friction. The tests were run on a laptop with an Intel® Core™ Ultra 7 165U.